5 log base 3 of 27
__________________
3 log base 3 of 81
It is written as a division problem, I couldn't get the line directly between the two equations.

Answer by oscargut(2103)

5 log base 3 of 27
__________________
3 log base 3 of 81
= 5 (3)
_____________
3 (4)
= 15/12
= 5/4
Answer: 5/4
If you need more help with math questions/assignments/tests
You can ask me more using this e-mail (mthman@gmail.com) or you can visit my site:
http://www.online-maths-help.com/
(Payment is using paypal)